OpenStudy (anonymous):

6.2 radians into degrees

OpenStudy (anonymous):

Use 180/π. 6.2(180/π) = ?

OpenStudy (anonymous):

For converting radians into degrees, we multiply 180/pi... where pi = 22/7 = 3.14

OpenStudy (anonymous):

355.23?

OpenStudy (anonymous):

Yeah approximately, you are right...

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

\[\frac{6.2}{\pi }180=355.234{}^{\circ} \]

OpenStudy (anonymous):

355.234